3.On six exams, Alex has scored 82, 86, 75, 91, 92, and 87 points out of 100. He wants to earn at least a B in the course, which corresponds to an overall average grade of 82 or better. Out of 100 points, what is the minimum number of points Alex must earn on the final exam to achieve this goal?
Answer by stanbon(48530)

On six exams, Alex has scored 82, 86, 75, 91, 92, and 87 points out of 100. He wants to earn at least a B in the course, which corresponds to an overall average grade of 82 or better. Out of 100 points, what is the minimum number of points Alex must earn on the final exam to achieve this goal?
--------------------------
Let his final exam score be "x":
Equation:
(82+86+75+91+92+87 + x )/7 >= 82
---
(513 + x)/7 >= 82
---
513 + x >= 574
x >= 61 (minimum score needed is 61)
=======================
Cheers,
Stan H.